### v1

# Conditional Generative Modeling for Enhanced Credit Risk Management in Supply Chain Finance

Abstract
The rapid expansion of cross-border e-commerce (CBEC) has created significant opportunities for small and medium-sized enterprises (SMEs), yet financing remains a critical challenge due to SMEs’ limited credit histories. Third-party logistics (3PL)-led supply chain finance (SCF) has emerged as a promising solution, leveraging in-transit inventory as collateral. We propose an advanced credit risk management framework tailored for 3PL-led SCF, addressing the dual challenges of credit risk assessment and loan size determination. Specifically, we leverage conditional generative modeling of sales distributions through Quantile-Regression-based Generative Metamodeling (QRGMM) as the foundation for risk estimation. We propose a unified framework that enables flexible estimation of multiple risk measures while introducing a functional risk measure formulation that systematically captures the relationship between these risk measures and varying loan levels, supported by theoretical guarantees. To capture complex covariate interactions in e-commerce sales data, we integrate QRGMM with Deep Factorization Machines (DeepFM). Extensive experiments on synthetic and real-world data validate the efficacy of our model for credit risk assessment and loan size determination. This study represents a pioneering application of generative AI in CBEC SCF risk management, offering a solid foundation for enhanced credit practices and improved SME access to capital.
keywords:
generative models , supply chain finance , credit risk management
[…]
[label1]organization=School of Management, Fudan University, city=Shanghai, postcode=200433, country=China
[…]
1 Introduction
E-commerce has become a cornerstone of global trade, with digital platforms enabling businesses of all sizes to reach international customers. In recent years, online retail has seen explosive growth worldwide: Business-to-Consumer (B2C) e-commerce sales have increased from approximately \$1.5 trillion in 2014 to over \$6 trillion in 2024, exceeding the GDPs of Japan and Germany (Huria 2019, World Bank 2025). Within this boom, cross-border e-commerce (CBEC) has emerged as a rapidly growing segment, increasing from around \$300 billion in 2015 to over \$1.2 trillion in 2024 (Future 2025). Despite this scale, access to financing remains a significant barrier—especially for small and medium-sized enterprises (SMEs). For instance, although over 70% of Amazon’s merchants source products internationally, only 1% obtain financial support from traditional banks (Jungle Scout 2024). These financing gaps are driven by SMEs’ limited collateral, thin credit histories, and the complexities of cross-border operations (Chiu 2024, Cai 2024).
Supply chain finance (SCF), particularly the third-party logistics (3PL)-led model, has emerged as a promising solution. In this model, 3PL providers collaborate with financial institutions to manage both logistics and financing, offering inventory-backed loans by leveraging real-time in-transit goods as collateral (Chen and Cai 2011, Ji et al. 2023). A salient example is UPS Capital, which combines logistics infrastructure with financial services (Huang et al. 2019).
Despite the promise of 3PL-led SCF, the practice of CBEC-oriented supply chain finance remains in its early stage, with fragmented adoption and limited standardization across markets. Industry developments reflect this nascent state: for instance, in 2024 the Hong Kong Export Credit Insurance Corporation just launched the market-first trade credit insurance programs specifically tailored to Hong Kong-based CBEC loans (Chiu 2024).This paper focuses on a specific 3PL-led CBEC-SCF mode where SMEs must pre-stock inventory in overseas warehouses to meet consumer expectations for rapid delivery, resulting in substantial working capital pressures.
In this 3PL-led CBEC-SCF mode, SMEs consign goods to 3PL-managed overseas warehouses, which serve as collateral for financial institutions. Lenders then assess credit risk based on the real-time value of in-transit inventory and predictive analytics of sales performance. The loan amount is adaptively adjusted: if sales forecasts indicate high volatility or low market demand, lenders may set loan-to-value ratios as low as 20% or reject financing entirely, while favorable forecasts may lead to higher loan amounts. This loan is secured against the products, which are managed within the 3PL’s coordinated logistics network. As the products eventually delivered, sales transactions on e-commerce platforms generate real-time cash flows. The revenue generated from these sales is then allocated primarily to repay the loan extended by the 3PL (with any surplus then remitted to the SME). This mode is akin to the CBEC-SCF modes proposed by Shi et al. (2023) and Ji et al. (2023), with the added innovation of lenders utilizing predictive sales analytics for credit risk control.
A central challenge in any lending model (and especially in CBEC-SCF) is effective risk management (Qiao and Zhao 2023). Two risk management questions stand out: (1) Credit risk assessment – how to reliably evaluate the probability of a borrower (SME) defaulting on a loan and estimate the potential financial loss and interested risk measures, and (2) Loan size determination – how to set credit limits or loan amounts in line with the borrower’s repayment capacity. These two questions are inherently interrelated and jointly shape the overall risk profile of lending decisions: the outcome of one directly affects the other. As such, addressing them in an integrated manner is essential for robust risk management.
In the 3PL-led CBEC-SCF model, where loan repayments depend on future sales revenues, accurate credit risk assessment and loan sizing hinge on modeling the conditional sales distribution $P(Y|\mathbf{x})$, given contextual features $\mathbf{x}$ such as product attributes (e.g., ratings, pricing, category), seller characteristics (e.g., review scores, platform tenure, fulfillment efficiency), and market factors (e.g., buyer preferences, seasonality). Traditional methods often provide only point estimates like the mean of sales $Y$ (Hastie et al. 2009, Mohri et al. 2018), whereas learning the full conditional distribution $P(Y|\mathbf{x})$ enables a nuanced view of sales variability, including dispersion and tail risk. This richer representation is key for assessing default probabilities and expected losses, as it captures how input features $\mathbf{x}$ influence the full range of potential sales outcomes $Y$, thereby supporting more informed and robust credit decisions.
We identify two promising and distinctive opportunities within this 3PL-led CBEC SCF mode. The first opportunity arises from the immense volume and richness of sales data from e-commerce platforms. Such data come with a variety of easily collectible features, offering a granular view of consumer behavior and sales trends. In contrast, existing credit risk assessment methods in SCF often center on the “SMEs itself-oriented” factors (e.g., credit histories, financial ratios, and operational metrics) and the “supply chain finance-oriented” factors (e.g., core enterprise credit ratings, industry trends, and supply chain concentration) to predict default risk via various classification models (Zhu et al. 2019, Hou et al. 2024, Belhadi et al. 2021). These conventional approaches face several limitations: the sample size is often limited (data about such diverse range of SCF influencing factors is frequently scattered across multiple entities, especially for SMEs); defaults are highly imbalanced (given the low incidence of bankruptcies), decreasing the accuracy of models (Hou et al. 2024); moreover, most of these models assume a fixed credit loan amount, lacking the flexibility required for adaptive loan sizing.
The second opportunity emerges from transformative advances in generative AI, which enable learning the entire conditional distribution critical for risk quantification. While early machine learning excelled at point prediction—estimating conditional expectations through supervised learning (Hastie et al. 2009, Mohri et al. 2018)—such approaches inherently obscure outcome variability. Generative models fundamentally redefine this paradigm. Pioneered by architectures like Generative Adversarial Networks (GANs, Goodfellow et al. 2014) and its conditional variations (CGAN, Mirza and Osindero 2014) and later advanced through Generative Pre-trained Transformers (GPTs, Radford et al. 2018), these generative models are designed to directly learn the entire conditional distribution of $Y$ given $\mathbf{x}$.
Given these opportunities, there is now a clear need for an accurate, robust, and efficient generative model that can learn the conditional distribution of the sales for specific products and covariates. Despite significant progress in generative AI—as highlighted by Hong et al. (2023)—most existing generative models are primarily designed for high-dimensional data (e.g., images, texts, audios, and videos), with an emphasis on generating outputs that are visually or semantically realistic rather than capturing the full underlying data distribution. In the context of SCF risk management, however, accurately and robustly acquiring the conditional distribution of low-dimensional variables like sales volume is critical. Poor forecasting can lead to critical misjudgments: overestimating the right tail of the distribution may result in a loan amount that exceeds the SME’s actual repayment capacity, heightening default risk, while underestimating sales potential might leave the SME under-financed, curtailing its growth and causing lenders to miss safe lending opportunities.
However, popular methods based on GAN structures, such as the CWGAN with gradient penalty (CWGAN-GP, Athey et al. 2021), often do not exhibit the desired level of accuracy and robustness in this setting (Hong et al. 2023). The current state-of-the-art technology for accurately and robustly capturing low-dimensional conditional distributions is Quantile-Regression-based Generative Metamodeling (QRGMM, Hong et al. 2023). In addition to its precision and robustness, QRGMM provides a significant advantage of efficiently generating a substantial number of random observations in real time, which is particularly valuable for estimating risk measures, facilitating the use of Monte Carlo methods to calculate various risk measures of interest, thereby enhancing our flexibility while minimizing estimation errors caused by sampling variance. In this paper, we explore the application of QRGMM for SCF credit risk management, aiming to leverage its capabilities to overcome the limitations inherent in traditional credit risk assessment methods.
While QRGMM offers a promising tool for learning the conditional distribution of sales volume, its application to SCF credit risk management is far from straightforward and presents several non-trivial challenges. Conventional SCF risk models are largely grounded in binary classification, estimating static default probabilities that fail to account for varying loan sizes or richer risk profiles. Even recent generative approaches using GANs focus on data augmentation, remaining confined to classification frameworks and failing to capture responsive loan size-risk relationships (Zhang et al. 2025). In real-world SCF settings, financial institutions demand adaptive tools that capture how risk metrics evolve with different loan levels, enabling real-time, data-driven loan sizing. Bridging generative modeling with practical SCF credit risk thus requires more than data augmentation: it demands a unified framework that translates the full conditional sales distribution output of QRGMM into actionable, loan-level-specific risk measures.
To meet this challenge, we develop an innovative functional formulation that effectively links the learned sales distribution to various credit risk assessments. Specifically, we introduce a unified framework that systematically formulates key risk metrics as explicit functions of the loan level: $r_{1}(l)$ for default probability, $r_{2}(l)$ for expected loss, and the versatile $r_{3}(l)$ for more generalized risk measures. By doing so, we provide a clear and efficient bridge between the generative sales volume model and a broad class of credit risk measures relevant to SCF. Importantly, this formulation accommodates a wide range of risk metrics within a unified and coherent structure, eliminating the need to construct and analyze cumbersome and inefficient separate static models for each risk measure individually, and thereby significantly enhancing decision-making flexibility. This flexibility allows practitioners to adaptively assess diverse credit risks across different loan sizes in real time, a capability that has been largely absent in prior methodologies.
Nevertheless, ensuring the reliability of this generative approach requires robust theoretical guarantees, which introduce another separate set of non-trivial challenges. While QRGMM ensures convergence in distribution for the estimated sales distribution, this alone is insufficient to guarantee the accurate estimation of more complex credit risk measures. In particular, key risk metrics like $r_{2}(l)$ and $r_{3}(l)$ involve non-linear, non-smooth transformations of the sales distribution, often expressed through complex expectations. Such structures prevent the direct application of standard convergence results from QRGMM, as ensuring accurate risk estimation requires uniform control not just over distribution estimates but also over the induced expectations across all loan levels. We overcome these difficulties by carefully exploiting the structures inherent to the SCF setting. Finally, we establish uniform convergence results for all formulated risk measures building upon the linear quantile regression model of Hong et al. (2023), providing a solid theoretical foundation for our unified framework, ensuring its utility and credibility for SCF credit risk management.
In practice, the above linear assumption seems to be rigid, potentially overlooking the complex relationships between product sales and their covariates in SCF. This complexity arises from the multifaceted nature of sales, which are influenced not only by the products and their attributes but also by various combinations of products, stores, and other factors affecting sales distributions. For example, in a cross-border e-commerce company that we worked with, product selection and sales strategies executed by diverse groups significantly affect outcomes due to varied skills and levels of collaboration. Traditional models, such as those used in recommendation systems like matrix factorization (MF, Lee and Seung 2000) and factorization machines (FM, Rendle 2010), primarily focus on predicting expected values rather than distributions. They are effective for learning combinatorial features but fall short in capturing the full range of the targeted distribution. To address this, we propose a novel approach that integrates QRGMM with deep factorization machines (DeepFM, Guo et al. 2017), enhancing accuracy in predicting target distributions. Extensive numerical experiments with both synthetic and real-world data validate our method’s ability to capture complex covariate interactions and precisely predict the conditional distribution of sales, which facilitates the estimation of pertinent credit risk measures and supports well-informed decision-making processes.
The remainder of the paper is structured as follows. Section 2 presents our functional risk formulation for loan-level determination in SCF, followed by Section 3 which introduces the quantile-regression-based generative model for SCF risk management and details its theoretical properties. Section 4 extends our framework by integrating DeepFM to effectively capture complex covariate interactions. Comprehensive numerical experiments with both synthetic and real-world data are presented in Section 5 to validate our approach. Finally, Section 6 concludes the paper and discusses potential avenues for future research.
2 Functional Risk Formulation for Loan-Level Determination in SCF
Consider an SME engaged in cross-border e-commerce that sells products via an online platform and requires 3PL-led supply chain financing. In this setting, the goods are managed by a third-party logistics provider during transit, and the revenue generated from eventual sales is primarily allocated to repay the loan. As a financial institution partnering with the 3PL, our goal is to determine an appropriate loan amount $l$ based on both the collateral value and projected sales performance, ensuring adequate risk control while aligning financing with the SME’s revenue potential.
Let $Y$ represent the random sales volume (in units) of the product over a fixed time horizon, let $r$ denote the net unit revenue per product sold, adjusted for shipping, handling, and platform operating costs. The effective revenue available for loan repayment is thus $rY$. We define the financial institution’s loss function as
$L(l)=(l-rY)^{+},$ (1)
where $(x)^{+}=\max\{x,0\}$. Intuitively, $L(l)$ captures the shortfall incurred when the loan amount $l$ exceeds the total revenue generated from sales, thereby encapsulating the risk associated with insufficient sales performance. This formulation abstracts the problem to its essential components, enabling us to evaluate how the loss (and hence the risk) varies as a function of the loan level.
The financial institution seeks to evaluate risk measures across the loan decision space $l\in[0,\bar{l}]$, where $\bar{l}$ denotes the maximum loan amount (often equal to the total collateral value to prevent arbitrage opportunities). We introduce a family of key risk measures that vary with $l$, offering flexibility in capturing different dimensions of credit risk. The first, the default probability, is defined as:
$r_{1}(l)=\Pr\{L(l)>0\}=\Pr\{l-rY>0\}=\Pr\{Y<l/r\},$
representing the likelihood that sales revenue fails to cover the loan. The second, the expected loss, is given by:
$r_{2}(l)=\mathbb{E}[L(l)]=\mathbb{E}[(l-rY)^{+}].$
quantifying the average shortfall in the event of default.
These two risk measures are fundamental to credit risk evaluation. For example, widely used credit risk measure, such as the Probability of Default (PD), is directly given by $r_{1}(l)$, and the Loss Given Default (LGD) can be computed as
$\text{LGD}=\frac{\mathbb{E}[L(l)\mid L(l)>0]}{l}=\frac{\mathbb{E}[L(l)]}{l\,\Pr\{L(l)>0\}}=\frac{r_{2}(l)}{l\,r_{1}(l)}.$
Moreover, suppose we are interested in the probability that the loss exceeds a threshold level $\xi$, where $0<\xi<l$. Then,
$\Pr\{L(l)>\xi\}$ $=\Pr\{(l-rS)^{+}>\xi\}$
$=\Pr\{l-rS>\xi\}$
$=\Pr\{S<(l-\xi)/r\}$
$=r_{1}(l-\xi).$
and the expected loss above threshold $\xi$ is given by
$\mathbb{E}\left[L(l)\,\mathbb{I}\{L(l)>\xi\}\right]=$ $~{}\mathbb{E}\left[(l-rY)\,\mathbb{I}\{l-rY>\xi\}\right]$
$=$ $~{}\mathbb{E}\left[(l-\xi-rY)\,\mathbb{I}\{rY<l-\xi\}\right]$
$+\xi\,\Pr\{rY<l-\xi\}$
$=$ $~{}r_{2}(l-\xi)+\xi\,r_{1}(l-\xi).$
This demonstrate that $r_{1}(l)$ and $r_{2}(l)$ serve as building blocks for broader risk evaluations.
In addition, to further broaden the scope of our framework, we can also propose a more generalized risk measure, $r_{3}(l)$, defined as:
$r_{3}(l)=\mathbb{E}\left[g_{l}(Y)\cdot\mathbb{I}\{Y<a(l)\}\right],$
where $g_{l}(\cdot)$ is a loan-level-dependent function that quantifies specific risk forms, and $a(l)$ is a threshold that varies with $l$, often tied to the default point (e.g., $a(l)=l/r$). This formulation further enhances flexibility, enabling $r_{3}(l)$ to adapt to a vastly wide range of risk metrics by tailoring $g_{l}(\cdot)$ and $a(l)$. The generality of $r_{3}(l)$ is most clearly evident in its ability to encompass both $r_{1}(l)$ and $r_{2}(l)$. By setting $g_{l}(Y)=1$ and $a(l)=l/r$, we obtain:
$r_{3}(l)=\mathbb{E}\left[1\cdot\mathbb{I}\{Y<l/r\}\right]=\Pr\{Y<l/r\}=r_{1}(l).$
Similarly, with $g_{l}(Y)=l-rY$ and $a(l)=l/r$, it follows that:
$r_{3}(l)=\mathbb{E}\left[(l-rY)\cdot\mathbb{I}\{Y<l/r\}\right]=\mathbb{E}[(l-rY)^{+}]=r_{2}(l).$
Beyond these foundational measures, $r_{3}(l)$ also extends to more complex risk assessments by allowing the loss function to take on more sophisticated, nonlinear forms. For instance, we can define $g_{l}(Y)=h(l-rY)$, where $h(x)=x^{p}$ for $p>0$ or other complex distortion functions. When $p=2$ and $a(l)=l/r$, this becomes $r_{3}(l)=\mathbb{E}[((l-rY)^{+})^{2}]$, which amplifies larger losses quadratically, emphasizing the severity of extreme shortfalls when sales fall well below the loan amount. Moreover, $g_{l}(Y)$ can be tailored to address scenario-dependent risks, such as those influenced by liquidity constraints or regulatory penalties by setting $g_{l}(Y)=(l-rY)\cdot(1+kY)$, where $k$ is an additional cost coefficient. Thus, through appropriate choices of $g_{l}(\cdot)$ and $a(l)$, $r_{3}(l)$ can offer a versatile framework for evaluating sophisticated risk measures, ensuring robustness and adaptability in diverse SCF applications and real-world financial challenges.
Now, the key challenge becomes efficiently computing these risk measures across all $l\in[0,\bar{l}]$. The crux of this challenge is to learn the distribution of $Y$. Suppose we have a generative model $\hat{Y}$ capable of producing $K$ independent and identically distributed observations $\{\hat{Y}_{i}\}_{i=1}^{K}$ rapidly, such that the distribution of $\hat{Y}$ closely approximates that of $Y$. Then, we can estimate the risk measures using Monte Carlo simulation (Hong et al. 2014):
$\hat{r}_{1,K}(l)=\frac{1}{K}\sum_{i=1}^{K}\mathbb{I}\{l-r\hat{Y}_{i}>0\},$
$\hat{r}_{2,K}(l)=\frac{1}{K}\sum_{i=1}^{K}\left(l-r\hat{Y}_{i}\right)^{+},$
$\hat{r}_{3,K}(l)=\frac{1}{K}\sum_{i=1}^{K}g_{l}(\hat{Y}_{i})\cdot\mathbb{I}\{\hat{Y}_{i}<a(l)\}.$ (2)
Therefore, the ability to generate these observations rapidly and accurately is crucial for evaluating risk measures across a continuum of loan levels. This motivates our use of a generative model to capture the full probability distribution of $Y$, enabling efficient and reliable risk assessments across all potential loan levels.
Our proposed functional risk assessment framework revolutionizes traditional SCF credit risk evaluations, which typically rely on methods like logistic regression or support vector machines to estimate a single risk measure—often the probability of default—for a fixed loan amount. By modeling risk measures such as $r_{1}(l)$, $r_{2}(l)$, and the versatile $r_{3}(l)$ as functions of loan size $l$, our functional approach provides a unified and comprehensive risk profile across a range of loan levels. Supported by a fast and reliable generative model and Monte Carlo approximations, this framework enables financial institutions to assess diverse credit risk measures and make informed loan sizing decisions, aligning financing with sales potential while significantly enhancing SCF risk management compared to static, single-point traditional methods.
3 Constructing a Conditional Generative Model for Sales Distribution
In this section, we first formulating the problem of constructing a conditional generative model for sale distribution. We then present QRGMM as an effective generative modeling technique that meets the specific requirements of SCF risk management. Finally, we establish theoretical guarantee for the learning of key risk measures, $r_{1}(l)$, $r_{2}(l)$, and $r_{3}(l)$, using QRGMM.
3.1 Problem Formulation
When a SME raises financing through 3PL-led CEBC-SCF mode, the sales revenue of its products will generally be prioritized to repay the loan. Hence, the random variable representing product sales, $Y(\mathbf{x})$, directly determines the loan repayment outcome. Here, the multi-dimensional vector $\mathbf{x}\in\mathcal{X}\subset\mathbb{R}^{p}$ denotes the associated covariates, such as product features, seller characteristics, and other auxiliary information. We collect $n$ historical observations $\{(\mathbf{x}_{i},y_{i})\}_{i=1}^{n}$ from e-commerce platforms, where $y_{i}$ are realizations of $Y(\mathbf{x}_{i})$. To effectively manage credit risk in SCF, our objective is to develop a generative model $\hat{Y}(\mathbf{x})$ utilizing the dataset $\{(\mathbf{x}_{i},y_{i})\}_{i=1}^{n}$. This model should accurately and robustly generate random observations that replicate the conditional distribution of $Y(\mathbf{x})$ given the covariates $\mathbf{x}$, while also ensuring rapid generation.
While generative models have significantly propelled advancements in areas such as image and video synthesis as well as language generation (Goodfellow et al. 2014, Radford et al. 2018, Ho et al. 2020, Liu et al. 2024), many of them are designed for high-dimensional data, focusing on visual or semantic realism rather than accurately capturing low-dimensional distributions, especially in the tails. Some models like GANs and variational autoencoders (Kingma and Welling 2013) also suffer from mode collapse issues (Saxena and Cao 2021), which reduces data diversity and fails to replicate the true distribution accurately, while others (e.g., diffusion-based approaches) are too slow for large-scale Monte Carlo simulation. In contrast, QRGMM (Hong et al. 2023) is designed to address these issues, offering a compelling alternative for SCF applications.
3.2 The QRGMM Algorithm
The foundation of QRGMM lies in the inverse transform method, where a uniform random variable $U$ on $(0,1)$ is mapped into a random observation from the target distribution by applying the conditional inverse cumulative distribution function (CDF) (also known as quantile function) $F_{Y}^{-1}(\tau|\mathbf{x})$. QRGMM approximates this inverse CDF by fitting quantile regressions on predefined grid points and then interpolating for intermediate quantiles. QRGMM operates in two stages: an offline learning stage where quantile regressions are fitted using the dataset $\{(\mathbf{x}_{i},y_{i})\}_{i=1}^{n}$ across a grid of quantile levels evenly spaced in the interval (0,1), and an online generation stage that generates uniform random variables and transforms them into target random variables by applying the fitted quantile function with linear interpolation. Algorithm 1 outlines the procedure.
Algorithm 1 QRGMM
Offline Stage:
  • Collect dataset $\{(\mathbf{x}_{i},y_{i})\}_{i=1}^{n}$.
  • Choose an integer $m$, and discretize $[0,1]$ by $\tau_{j}=\tfrac{j}{m},\;j=1,\ldots,m-1$.
  • Fit quantile regression models $\hat{Q}_{\tau_{j}}(\mathbf{x})$ for all $j=1,\ldots,m-1$.
Online Stage:
  • Given new covariates $\mathbf{x}$.
  • Generate $u_{1},\ldots,u_{K}$ i.i.d. from $\mathrm{Uniform}(0,1)$.
  • For $k=1,\ldots,K$, output $\hat{Y}_{k}(\mathbf{x})$ as follows:
    for $u_{k}\in[\tau_{j},\tau_{j+1})$ with $j=1,\ldots,m{-}2$,
    $\hat{Y}_{k}(\mathbf{x})=\hat{Q}_{\tau_{j}}(\mathbf{x})+m(u_{k}-\tau_{j})\left[\hat{Q}_{\tau_{j+1}}(\mathbf{x})-\hat{Q}_{\tau_{j}}(\mathbf{x})\right].$
    otherwise,
    $\hat{Y}_{k}(\mathbf{x})=\begin{cases}\hat{Q}_{\tau_{1}}(\mathbf{x}),&u_{k}<\tau_{1},\\ \hat{Q}_{\tau_{m-1}}(\mathbf{x}),&u_{k}\geq\tau_{m-1}.\end{cases}$
Why QRGMM for SCF?
Below, we elaborate on the key features of QRGMM in detail, highlighting why it is a compelling choice for risk management in SCF.
  • Accuracy and Robustness: By explicitly modeling conditional quantiles at multiple levels, QRGMM learns the full shape of a distribution rather than focusing merely on mean predictions or specific performance measures. This quantile-based approach inherently avoids issues such as mode collapse, and it provides more reliable estimates of the tails of the distribution which have more impact on the SCF credit risk. Unlike adversarial models such as GANs that require solving challenging saddle-point optimization (notoriously difficult to tune), QRGMM merely involves fitting quantile regressions and selecting the main parameter—the number of grid points $m$. Furthermore, the choice of $m=O(\sqrt{n})$, guided by theoretical and empirical insights of Hong et al. (2023), effectively balances performance with computational cost, resulting in a simpler and more stable training process.
  • Speed: Numerical results in Hong et al. (2023) show that QRGMM can generate a lot of observations in an extremely short time—e.g., 100,000 random samples in under 0.01 seconds. This efficiency stems from its reliance on simple linear interpolation in the online generation stage, after the quantile regressions have been fitted offline. Such rapid sampling is particularly valuable in SCF risk management, as it enables real-time generation of extensive samples, significantly reducing the computational cost of sample generation. This allows us to produce large sample sets to support Monte Carlo methods, computing various risk measures of interest with enhanced flexibility, while minimizing estimation errors caused by sampling variance.
In summary, by framing SCF risk management through the lens of the full conditional distribution of product sales, we underscore the need for a generative model that is not only accurate and robust but also computationally efficient. QRGMM effectively meets these requirements, providing a promising tool for the integrated framework that links credit risk evaluation with loan sizing decisions. However, to fully realize the practical potential of QRGMM in SCF applications, it is essential to go beyond its generative capabilities and rigorously establish how its learning outcomes translate into reliable credit risk measures. Specifically, building on the problem formulation in Section 2, the next critical step is to provide strong theoretical guarantees on the estimation of the key risk measures—namely, $r_{1}(l)$, $r_{2}(l)$, and $r_{3}(l)$—across the entire range of loan levels $l\in[0,\bar{l}]$. This ensures that the QRGMM-based framework is not only practically effective but also theoretically sound, setting the stage for robust real-world deployment.
3.3 Asymptotic Analysis of Risk Measure Estimations via QRGMM
In this subsection, we establish theoretical guarantees for the estimation of $r_{1}(l)$, $r_{2}(l)$, and $r_{3}(l)$ for all $l\in[0,\bar{l}]$, thereby ensuring asymptotic validity for a broad range of SCF risk measure estimations through QRGMM. Prior to proceeding, we emphasize several key considerations.
Model Assumption
In this section, we adopt the linear quantile regression model assumption of Hong et al. (2023), which lay the theoretical groundwork for analyzing the asymptotic properties of QRGMM. Suppose $Y(\mathbf{x})$ is a continuous random variable on its support given the input $\mathbf{x}$. Its conditional distribution and density functions are denoted as $F_{Y}(\cdot|\mathbf{x})$ and $f_{Y}(\cdot|\mathbf{x})$. The linear quantile regression model assumption is stated as follows.
Assumption 1.
$F^{-1}_{Y}(\tau|\mathbf{x})=\boldsymbol{\beta}(\tau)^{\intercal}\mathbf{x}$ all $\tau\in(0,1)$ and $\boldsymbol{\beta}(\tau)$ is continuous in $\tau$.
Assumption 1 posits that the linear quantile regression model accurately represents the true model of the conditional quantile functions of $Y$ given $\mathbf{x}$ for all $\tau$ in $(0,1)$, implying that $F^{-1}_{Y}(\tau|\mathbf{x})$ is linear with respect to the covariates $\mathbf{x}$. Although Assumption 1 may seem restrictive, it is widely used in practice due to its simplicity and interpretability. This assumption is foundational not only because it simplifies the theoretical framework but also because it is underpinned by a substantial literature on classical quantile regression, which provides the necessary theoretical foundation for establishing the asymptotic analysis. Furthermore, in conjunction with domain knowledge, it can be expanded to capture nonlinearity by substituting $\mathbf{x}$ with a set of nonlinear basis functions $\boldsymbol{b(\mathbf{x})}$. In terms of theoretical analysis, the theoretical results can be readily extended when employing basis functions. Supporting evidence from numerical experiments by Hong et al. (2023) demonstrates that even when the linear model assumption does not strictly hold, the approximation using basis functions still produces favorable outcomes. In Section 4, we will explore an extension of the linear quantile regression model to the DeepFM model, enhancing its ability to capture complex combinatory features.
Under Assumption 1, the quantile function model $\hat{Q}_{\tau_{j}}(\mathbf{x})$ in Algorithm 1 can then be obtained by $\hat{Q}_{\tau_{j}}(\mathbf{x})=\hat{\boldsymbol{\beta}}({\tau_{j}})^{\intercal}\mathbf{x}$ for all $j=1,\ldots,m-1$, where $\hat{\boldsymbol{\beta}}(\tau_{j})$ is the estimation of $\boldsymbol{\beta}(\tau_{j})$ using the linear quantile regression on the dataset $\{(\mathbf{x}_{i},y_{i})\}_{i=1}^{n}$ such that
$\hat{\boldsymbol{\beta}}({\tau_{j}})=\underset{\boldsymbol{\beta}}{\text{argmin}}\sum_{i=1}^{n}\rho_{\tau_{j}}(y_{i}-\boldsymbol{\beta}^{\intercal}\mathbf{x}_{i}),\quad j=1,2,\ldots,m-1.$
where $\rho_{\tau}(u)=\left[\tau-I(u\leq 0)\right]u$ is the pinball loss function of quantile regression, this loss penalizes over- and under-predictions asymmetrically, guiding the model to estimate the desired quantile. And the above optimization problem can be efficiently solved via linear programming (Koenker 2005).
Source of Error
Our credit risk estimation procedure builds upon QRGMM to produce a generative model $\hat{Y}(\mathbf{x})$ that approximates the distribution of $Y(\mathbf{x})$. Given $\hat{Y}(\mathbf{x})$, we then rely on Monte Carlo simulation to estimate risk measures by generating $K$ samples $\{\hat{Y}_{k}(\mathbf{x})\}_{k=1}^{K}$ from the generative model and transforming them into the estimation of risk measures through (2). This process involves three distinct sources of approximation:
  1. 1.
    Quantile Estimation Error: The construction of $\hat{Y}(\mathbf{x})$ begins with using quantile regression to estimate a discrete set of conditional quantile functions $\hat{Q}_{\tau_{j}}(\mathbf{x})$ from an offline dataset of size $n$. These finite-sample quantile function estimates at grid points inherently deviate from the true quantile functions due to data variability and model assumptions.
  2. 2.
    Interpolation Error: To approximate $F^{-1}_{Y}(\tau|\mathbf{x})$ for all $\tau$, we rely on a set of $m$ quantile levels and piecewise interpolation. This discretization, governed by $m$, introduces interpolation error that diminishes as we increase the number of quantile levels.
  3. 3.
    Monte Carlo Sampling Error: After constructing $\hat{Y}(\mathbf{x})$, the final estimates of risk measures are computed by sampling from this approximate distribution. The finite number $K$ of Monte Carlo samples leads to sampling variability, which decreases as $K$ grows.
Notice that the third source of error can be made arbitrarily small by choosing a sufficiently large $K$, which is computationally feasible since QRGMM enables extremely fast data generation. Therefore, we focus our theoretical analysis on the first two error sources: the quantile estimation error associated with the finite sample size $n$, and the interpolation error linked to the number of quantile levels $m$.
3.3.1 Asymptotic Analysis for $r_{1}(l)$ Estimation
In this subsection, we establish the asymptotic properties of the estimated default probability $\hat{r}_{1}(l)$, a key metric in credit risk assessment for SCF. Given any $\mathbf{x}$, the default probability $r_{1}(l)$ satisfies:
$r_{1}(l)=\Pr\{L(l)>0\}=\Pr\{Y(\mathbf{x})<l/r\}=F_{Y}\bigl{(}\,l/r\!\mid\!\mathbf{x}\bigr{)},$
In the QRGMM framework, ignoring the Monte Carlo sampling error, $r_{1}(l)$ is essentially approximated by $\hat{r}_{1}(l)=F_{\hat{Y}}(\,l/r|\mathbf{x})$, where $F_{\hat{Y}}(\cdot|\mathbf{x})$ is the CDF of the generative model $\hat{Y}(\mathbf{x})$. The accuracy of $\hat{r}_{1}(l)$ thus hinges on the convergence of $F_{\hat{Y}}(y|\mathbf{x})$ to the true CDF $F_{Y}(y|\mathbf{x})$.
In SCF context, by understanding the distribution of $Y(\mathbf{x})$, lenders can determine the appropriate loan level $l$ such that we can achieve a desired PD. To support this loan sizing decision making, it is not enough to have mere pointwise convergence of $F_{\hat{Y}}(\,l/r|\mathbf{x})$ for a fixed $l$. Instead, a stronger form of uniform convergence is required. Uniform convergence of $F_{\hat{Y}}(y|\mathbf{x})$ for all $y\in[0,\infty)$ ensures the uniform convergence of $\hat{r}_{1}(l)$ for all $l\in[0,\bar{l}]$. This uniform convergence is a stronger and more challenging requirement, as it ensures consistent accuracy across the entire loan decision space, addressing a gap in existing literature that often overlooks the responsive interplay between loan levels and risk measures.
Our objective is to prove both pointwise and uniform convergence of $\hat{r}_{1}(l)$ to $r_{1}(l)$. These theoretical guarantees are essential for ensuring the reliability of QRGMM in SCF. Leveraging results from Hong et al. (2023), we proceed to establish the uniform convergence of $F_{\hat{Y}}(y|\mathbf{x})$, which in turn guarantees the uniform convergence of $\hat{r}_{1}(l)$ over the loan decision space.
The Uniform Convergence of $F_{\hat{Y}}(y|\mathbf{x})$
We begin by recalling their assumptions and results.
Assumption 2.
  1. (2.a)
    $F_{Y}(\cdot|\mathbf{x})$ is absolutely continuous on a common support for all $\mathbf{x}\in\mathcal{X}$.
  2. (2.b)
    For any given interval $[\tau_{\ell},\tau_{u}]\subset(0,1)$ with $10<\tau_{\ell}<\tau_{u}<$ 1, the density functions $f_{Y}(\cdot|\mathbf{x})$ are uniformly bounded away from 00 and $\infty$ in the sense that
    $0<\inf_{\tau\in[\tau_{\ell},\tau_{u}]}\inf_{\mathbf{x}\in\mathcal{X}}f_{Y}(F_{Y}^{-1}(\tau|\mathbf{x})|\mathbf{x})<\infty,$
    $0<\sup_{\tau\in[\tau_{\ell},\tau_{u}]}\sup_{\mathbf{x}\in\mathcal{X}}f_{Y}(F_{Y}^{-1}(\tau|\mathbf{x})|\mathbf{x})<\infty.$
Assumption 3.
  1. (3.a)
    For any given interval $[\tau_{\ell},\tau_{u}]\subset(0,1)$ with $10<\tau_{\ell}<\tau_{u}<$ 1, there exist positive definite matrices $D_{0}$ and $D_{1}(\tau)$ such that
    $\lim_{n\to\infty}\frac{1}{n}\sum_{i=1}^{n}\mathbf{x}_{i}\mathbf{x}_{i}^{\intercal}=D_{0},$
    $\lim_{n\to\infty}\frac{1}{n}\sum_{i=1}^{n}f_{Y}(F^{-1}_{Y}(\tau|\mathbf{x}_{i})|\mathbf{x}_{i})\mathbf{x}_{i}\mathbf{x}_{i}^{\intercal}=D_{1}(\tau)$
    uniformly for $\tau\in[\tau_{\ell},\tau_{u}].$
  2. (3.b)
    $\max_{1\leq i\leq n}\|\mathbf{x}_{i}\|=o(\sqrt{n})$.
Assumption 2 and Assumption 3 are commonly adopted in quantile regression analyses to ensure robust and reliable statistical inferences. These conditions facilitate the estimation of quantile regression parameters and support the theoretical underpinnings of asymptotic analyses.
With the above assumptions, the following uniform convergence of $F_{\hat{Y}}(y|\mathbf{x})$ can be obtained.
Lemma 1.
Under Assumptions 13, the following results hold:
  1. (1.a)
    For any $\mathbf{x}\in\mathcal{X}$, $\hat{Y}(\mathbf{x})\Rightarrow Y(\mathbf{x})$ as $n,m\to\infty$.
  2. (1.b)
    For any $\mathbf{x}\in\mathcal{X}$, $\sup\limits_{y\in[0,\infty)}\left|F_{\hat{Y}}(y|\mathbf{x})-F_{Y}(y|\mathbf{x})\right|\to 0$ as $n,m\to\infty$.
where $\Rightarrow$” denotes convergence in distribution.
Asymptotics for $r_{1}(l)$ Estimation
Recall that $\hat{r}_{1}(l)$ and $r_{1}(l)$ represent the estimated and true probabilities of default, respectively. Their difference is precisely:
$\bigl{|}\hat{r}_{1}(l)-r_{1}(l)\bigr{|}\;=\;\Bigl{|}\,F_{\hat{Y}}\bigl{(}\,l/r\!\mid\!\mathbf{x}\bigr{)}\;-\;F_{Y}\bigl{(}\,l/r\!\mid\!\mathbf{x}\bigr{)}\Bigr{|}.$
Hence, by Lemma 1, we obtain:
Theorem 1.
Under Assumptions 13, the following results hold:
  1. (1.a)
    For any $\mathbf{x}\in\mathcal{X}$, and for any $l\in[0,\bar{l}]$, we have $\hat{r}_{1}(l)\to r_{1}(l)$ as $n,m\to\infty$.
  2. (1.b)
    For any $\mathbf{x}\in\mathcal{X}$, we have $\sup\limits_{l\in[0,\bar{l}]}\left|\hat{r}_{1}(l)-r_{1}(l)\right|\to 0$ as $n,m\to\infty$.
These results ensure $\hat{r}_{1}(l)$’s uniform convergence across $l\in[0,\bar{l}]$, supporting adaptive loan sizing in SCF. This analysis not only offers a rigorous foundation for practitioners to trust QRGMM-generated PD estimation but also lays the groundwork for analyzing more complex risk measures like $r_{2}(l)$ and $r_{3}(l)$, explored in subsequent subsections.
3.3.2 Asymptotic Analysis for $r_{2}(l)$ Estimation
In this subsection, we delve deeper into the analysis of $r_{2}(l)=\mathbb{E}\left[(l-rY)^{+}\right],$ and its approximation via QRGMM. Compared to the estimation of the default probability $r_{1}(l)$, which involves the direct evaluation of the estimated conditional distribution function $F_{\hat{Y}}(y|\mathbf{x})$, the estimation of $r_{2}(l)$ is more subtle. It requires analyzing the convergence of the expectation of a transformation of the random variable $Y(\mathbf{x})$, namely the positive part $(l-rY)^{+}$. This complicates the asymptotic analysis since standard weak convergence results do not automatically guarantee convergence of expectations for general functions.
Fortunately, the structure of $r_{2}(l)$ offers an advantage: for all $l\in[0,\bar{l}]$, the function $y\mapsto(l-ry)^{+}$ is uniformly bounded by $\bar{l}$ and continuous in $y$. These properties enable us to invoke a classical result from probability theory regarding the convergence of expectations under weak convergence:
Lemma 2 (Portmanteau Theorem).
Let $Z_{n}\Rightarrow Z$ as $n\to\infty$. Then for any bounded and continuous function $h$, it holds that $\mathbb{E}\left[h(Z_{n})\right]\to\mathbb{E}\left[h(Z)\right].$
This Portmanteau theorem (see, e.g., Van der Vaart (2000, page 6)) provides a critical tool in establishing uniform convergence of $\hat{r}_{2}(l)$. Specifically, for each $l\in[0,\bar{l}]$, we define the transformation $f_{l}(y):=(l-ry)^{+},$ which is uniformly bounded above by $\bar{l}$ and continuous on $[0,\infty)$. Given that $\hat{Y}(\mathbf{x})\Rightarrow Y(\mathbf{x})$ as $n,m\to\infty$ by Lemma (1.a), applying Lemma 2 ensures pointwise convergence $\hat{r}_{2}(l)\to r_{2}(l)$ for each $l$. Furthermore, leveraging the equicontinuity of $f_{l}(y)=(l-ry)^{+}$ with respect to $l$, we can extend this result to uniform convergence over $l\in[0,\bar{l}]$. The result is summarized in the following theorem.
Theorem 2.
Suppose that Assumptions 13 hold, if $m$ satisfies $\frac{1}{m}\to 0$ and $\frac{n}{m}\to\infty$ as $n\to\infty$, then, for any given $\mathbf{x}$, as $m,n\to\infty$, we have
$\sup\limits_{l\in[0,\bar{l}]}\bigl{|}\hat{r}_{2}(l)-r_{2}(l)\bigr{|}\to 0.$
Proof.
For any $l\in[0,\bar{l}]$, the function $f_{l}(y)=(l-ry)^{+}$ is Lipschitz continous in $y$ with Lipschitz constant $r$, and uniformly bounded with $0\leq f_{l}(y)\leq\bar{l}$ for all $y\in\mathbb{R}$. Additionally, for all $y\in\mathbb{R}$ and $l,l^{\prime}\in[0,\bar{l}]$, we have
$|f_{l}(y)-f_{l^{\prime}}(y)|\leq|l-l^{\prime}|,$
which implies that $\{f_{l}\}$ is equicontinuous in the parameter $l$ (since each $f_{l}(y)$ is Lipschitz continuous in $l$ with a common Lipschitz constant constant $1$ regardless of $y$).
For any $\mathbf{x}\in\mathcal{X}$, since $\hat{Y}(\mathbf{x})\Rightarrow Y(\mathbf{x})$ as $n,m\to\infty$ by Lemma (1.a) and each $f_{l}$ is continuous and bounded, by Lemma 2, it follows that
$\hat{r}_{2}(l)=\mathbb{E}[f_{l}(\hat{Y}(\mathbf{x}))]\to r_{2}(l)=\mathbb{E}[f_{l}(Y(\mathbf{x}))]$
for every fixed $l\in[0,\bar{l}]$. To strengthen this to uniform convergence over $l\in[0,\bar{l}]$, fix any $\epsilon>0$. By the compactness of $[0,\bar{l}]$ and the equicontinuity of $\{f_{l}\}$ in $l$, there exists a finite $\delta$-net $\{l_{1},\dots,l_{T}\}\subset[0,\bar{l}]$ such that for any $l\in[0,\bar{l}]$, there exists $j\in\{1,\dots,T\}$ with $|l-l_{j}|<\delta$, where $\delta:=\epsilon/3$.
Then for such $l$ and corresponding $l_{j}$, we decompose the total error as:
$\left|\mathbb{E}[f_{l}(\hat{Y}_{n})]-\mathbb{E}[f_{l}(Y)]\right|$
$\leq$ $\left|\mathbb{E}[f_{l}(\hat{Y}_{n})]-\mathbb{E}[f_{l_{j}}(\hat{Y}_{n})]\right|$
$+$ $\left|\mathbb{E}[f_{l}(\hat{Y}_{n})]-\mathbb{E}[f_{l_{j}}(\hat{Y}_{n})]\right|$
$+$ $\left|\mathbb{E}[f_{l_{j}}(Y)]-\mathbb{E}[f_{l}(Y)]\right|.$
The first and third terms are bounded by
$\left|\mathbb{E}[f_{l}(\hat{Y}_{n})]-\mathbb{E}[f_{l_{j}}(\hat{Y}_{n})]\right|$
$=$ $\left|\mathbb{E}[f_{l}(\hat{Y}_{n})-f_{l_{j}}(\hat{Y}_{n})]\right|$
$\leq$ $\left|\mathbb{E}[l-l_{j}]\right|=|l-l_{j}|<\delta=\epsilon/3,$
and similarly for $\left|\mathbb{E}[f_{l_{j}}(Y)]-\mathbb{E}[f_{l}(Y)]\right|$.
For the second term, since the set $\{l_{1},\dots,l_{T}\}$ is finite and $\mathbb{E}[f_{l_{j}}(\hat{Y}_{n})]\to\mathbb{E}[f_{l_{j}}(Y)]$ for each $j$, there exists $M$ and $N$ such that for all $m>M$, $n>N$ and all $j$,
$\left|\mathbb{E}[f_{l_{j}}(\hat{Y}_{n})]-\mathbb{E}[f_{l_{j}}(Y)]\right|<\epsilon/3.$
Combining the above, we conclude that for all $l\in[0,\bar{l}]$ and all $m,n$ sufficiently large,
$\left|\mathbb{E}[f_{l}(\hat{Y}_{n})]-\mathbb{E}[f_{l}(Y)]\right|<\epsilon.$
Therefore,
$\sup\limits_{l\in[0,\bar{l}]}\bigl{|}\hat{r}_{2}(l)-r_{2}(l)\bigr{|}=\sup_{l\in[0,\bar{l}]}\left|\mathbb{E}[f_{l}(\hat{Y}_{n})]-\mathbb{E}[f_{l}(Y)]\right|\to 0.$
This asymptotic guarantee is particularly significant in the SCF setting, where the continous and bounded structure of $f_{l}(y)$ simplify the analysis. Without these properties, expectation and tail behavior could complicate the convergence analysis, requiring stronger assumptions. Our approach effectively exploits this structure, leading to the uniform convergence for $r_{2}(l)$ estimation. This result provides theoretical validation for the use of QRGMM in estimating expected loss risk measures, reinforcing its applicability in SCF risk management.
3.3.3 Asymptotic Analysis for $r_{3}(l)$ Estimation
Here, ignoring the MC error, we prove the uniform convergence of the estimated risk measure $\hat{r}_{3}(l)=\mathbb{E}[g_{l}(\hat{Y}(\mathbf{x}))\cdot\mathbf{I}_{\{\hat{Y}(\mathbf{x})<a(l)\}}]$, computed using the QRGMM-based generative model $\hat{Y}(\mathbf{x})$, to the true $r_{3}(l)$. Unlike $r_{2}(l)$, where the function $(l-ry)^{+}$ is continuous and bounded, enabling the use of the Portmanteau Theorem to establish convergence, the general form of $g_{l}(y)\cdot\mathbf{I}_{\{y<a(l)\}}$ poses a challenge. The indicator function may introduce the discontinuity, rendering the transformation non-continuous. Thus, theoretical analysis method applied to $\hat{r}_{2}(l)$ is not directly applicable to $\hat{r}_{3}(l)$
To address this, we take a different approach to establish the uniform convergence for $\hat{r}_{3}(l)$, which is consequently also applicable to $\hat{r}_{1}(l)$ and $\hat{r}_{2}(l)$. Specifically, we first impose the following regularity conditions on $g_{l}(\cdot)$ and $a(l)$:
Assumption 4.
  1. (4.a)
    $a(l)$ is monotone on $[0,\bar{l}]$, satisfies $0\leq a(l)\leq a(\bar{l})<\infty$, with $a(l)\to 0$ as $l\to 0$.
  2. (4.b)
    $g_{l}(\cdot)$ is differentiable and Lipschitz continuous with constant $K$ within $y\in[0,a(\bar{l})]$ uniformly over $l\in[0,\bar{l}]$, and uniformly bounded by $g_{\bar{l}}(0)<\infty$.
Assumption (4.a) requires $a(l)$, a threshold function, to be monotonic in $l$ and and vanish as $l\to 0$, reflecting the intuition that smaller loans imply lower trigger levels. Assumption (4.b) ensures that $g_{l}(\cdot)$, representing loss in SCF, is differentiable, Lipschitz continuous with constant $K$ (implying $|g_{l}^{\prime}(y)|\leq K$) within $y\in[0,a(\bar{l})]$, and bounded by $M=g_{\bar{l}}(0)$, reflecting the maximum loss in the worst case that sales $Y=0$ and loan $l=\bar{l}$. These conditions, intuitive and reasonable for SCF risk management, ensure tractable analysis of $\hat{r}_{3}(l)$, leading to the following result.
Theorem 3.
Under Assumptions 14, if $m$ satisfies $\frac{1}{m}\to 0$ and $\frac{n}{m}\to\infty$ as $n\to\infty$, then for any given $\mathbf{x}$, as $n,m\to\infty$, we have
$\sup_{l\in[0,\bar{l}]}\left|\hat{r}_{3}(l)-r_{3}(l)\right|\to 0,$
Proof.
Recall that
$r_{3}(l)=\mathbb{E}\left[g_{l}(Y(\mathbf{x}))\cdot\mathbf{I}_{\{Y(\mathbf{x})<a(l)\}}\right],$
$\hat{r}_{3}(l)=\mathbb{E}\left[g_{l}(\hat{Y}(\mathbf{x}))\cdot\mathbf{I}_{\{\hat{Y}(\mathbf{x})<a(l)\}}\right].$
Express the expectations:
[…]
The difference is:
$|\hat{r}_{3}(l)-r_{3}(l)|=\left|\int_{0}^{a(l)}g_{l}(y)\,d(F_{\hat{Y}}-F_{Y})(y|\mathbf{x})\right|.$
As $g_{l}(y)$ is differentiable and Lipschitz continuous with constant $K$ within $y\in[0,a(\bar{l})]$ uniformly over $l\in[0,\bar{l}]$, we have $|g_{l}^{\prime}(y)|\leq K$ within $y\in[0,a(\bar{l})]$. Let
$\delta_{n,m}=\sup_{y\geq 0}|F_{\hat{Y}}(y|\mathbf{x})-F_{Y}(y|\mathbf{x})|,$
and denote $M=g_{\bar{l}}(0)$.
For any $\mathbf{x}\in\mathcal{X}$ and $l\in[0,\bar{l}]$, compute the difference:
$|\hat{r}_{3}(l)-r_{3}(l)|$
$=$ $\left|\int_{0}^{a(l)}g_{l}(y)\,dF_{\hat{Y}}(y|\mathbf{x})-\int_{0}^{a(l)}g_{l}(y)\,dF_{Y}(y|\mathbf{x})\right|$
$=$ $\left|\int_{0}^{a(l)}g_{l}(y)\,d(F_{\hat{Y}}-F_{Y})(y|\mathbf{x})\right|.$
Using integration by parts:
$\int_{0}^{a(l)}g_{l}(y)\,d(F_{\hat{Y}}-F_{Y})(y|\mathbf{x})$
$=$ $\left[g_{l}(y)\left(F_{\hat{Y}}(y|\mathbf{x})-F_{Y}(y|\mathbf{x})\right)\right]\big{|}_{0}^{a(l)}$
[…]
Evaluate the boundary term:
$\left[g_{l}(y)\left(F_{\hat{Y}}(y|\mathbf{x})-F_{Y}(y|\mathbf{x})\right)\right]\big{|}_{0}^{a(l)}$
$=$ $g_{l}(a(l))\left(F_{\hat{Y}}(a(l)|\mathbf{x})-F_{Y}(a(l)|\mathbf{x})\right)$
-- $g_{l}(0)\left(F_{\hat{Y}}(0|\mathbf{x})-F_{Y}(0|\mathbf{x})\right).$
Thus:
$\hat{r}_{3}(l)-r_{3}(l)$
$=$ $g_{l}(a(l))\left(F_{\hat{Y}}(a(l)|\mathbf{x})-F_{Y}(a(l)|\mathbf{x})\right)$
-- $g_{l}(0)\left(F_{\hat{Y}}(0|\mathbf{x})-F_{Y}(0|\mathbf{x})\right)$
[…]
For the boundary terms, since $|g_{l}(a(l))|\leq M$, $|g_{l}(0)|\leq M$, and $|F_{\hat{Y}}(y|\mathbf{x})-F_{Y}(y|\mathbf{x})|\leq\delta_{n,m}$, we have
$|g_{l}(a(l))|\cdot|F_{\hat{Y}}(a(l)|\mathbf{x})-F_{Y}(a(l)|\mathbf{x})|\leq M\delta_{n,m},$
$|g_{l}(0)|\cdot|F_{\hat{Y}}(0|\mathbf{x})-F_{Y}(0|\mathbf{x})|\leq M\delta_{n,m}.$
For the integral term, since $|g_{l}^{\prime}(y)|\leq K$ and $|F_{\hat{Y}}(y|\mathbf{x})-F_{Y}(y|\mathbf{x})|\leq\delta_{n,m}$, we have
$\int_{0}^{a(l)}|F_{\hat{Y}}(y|\mathbf{x})-F_{Y}(y|\mathbf{x})|\cdot|g_{l}^{\prime}(y)|\,dy$
[…]
as $a(l)\leq a(\bar{l})$. Combine them, we have
$|\hat{r}_{3}(l)-r_{3}(l)|$
$\leq$ $M\delta_{n,m}+M\delta_{n,m}+K\delta_{n,m}a(\bar{l})$
$=$ $\left(2M+Ka(\bar{l})\right)\delta_{n,m}.$
Since the bound is independent of $l$, it holds uniformly:
$\sup_{l\in[0,\bar{l}]}|\hat{r}_{3}(l)-r_{3}(l)|\leq(2M+Ka(\bar{l}))\delta_{n,m}.$
By Lemma 1(b), we have $\delta_{n,m}\to 0$, as $n,m\to\infty$. Thus:
$\sup_{l\in[0,\bar{l}]}|\hat{r}_{3}(l)-r_{3}(l)|\to 0.$
This completes the proof. ∎
This theorem confirms the uniform consistency of the QRGMM estimator for $r_{3}(l)$. The convergence guarantee is facilitated by the boundedness and smoothness of the integrand $g_{l}(\cdot)$, and the controlled behavior of the truncation threshold $a(l)$. These structural conditions are naturally satisfied in SCF applications,
The uniform convergence of $\hat{r}_{3}(l)$, matching the asymptotic behavior of $\hat{r}_{1}(l)$ and $\hat{r}_{2}(l)$, underscores the robustness of QRGMM in the SCF context. This theorem, coupled with the flexibility of $g_{l}(\cdot)$ and $a(l)$, enhancing the applicability of $r_{3}(l)$, allows $r_{3}(l)$ to adapt to diverse risk management needs while maintaining estimation consistency. In summary, the uniform convergence results of $\hat{r}_{1}(l)$, $\hat{r}_{2}(l)$, and $\hat{r}_{3}(l)$ ensures reliable risk assessment across varying loan levels, enhancing decision-making in SCF applications such as adaptive loan sizing and risk mitigation.
4 Extension of QRGMM for Practical Applicability in SCF Risk Management
In the previous chapters, we focused on the QRGMM under relatively simple model assumptions, notably linear quantile regression. While these assumptions facilitated theoretical derivations and provided valuable insights into convergence properties, they may be restrictive for real-world SCF scenarios. In practice, the relationships governing sales—especially in cross-border e-commerce—are often complex, involving intricate interactions among products, stores, and other contextual factors. Linear models may struggle to capture such combinational effects, potentially undermining the accuracy and effectiveness of distributional predictions required for robust SCF risk management.
To overcome these limitations, this chapter integrates advanced techniques from recommendation systems, specifically Deep Factorization Machines (DeepFM, Guo et al. 2017), with the QRGMM framework. This integration allows us to model non-linear, intricate feature interactions and improve the realism and accuracy of generated sales distributions. We term this extended model as QRGMM, highlighting its advanced capability to handle complex combinational features and non-linear relationships, essential for robust SCF applications.
4.1 Overview of Related Techniques
This subsection introduces the key building blocks for our extended model. We first discuss Matrix Factorization (MF, Lee and Seung 2000) and Factorization Machines (FM, Rendle 2010), fundamental techniques in recommendation systems that enable efficient modeling of latent and combinational features. We then present Deep Factorization Machines, which combine FM with deep learning technique to capture second-order and higher-order feature interactions. Lastly, the Quantile Regression Neural Network (QRNN, Cannon 2011) provides a remarkable approach for extending quantile regression from linear models to neural networks, making it an ideal glue to seamlessly integrate DeepFM with the QRGMM framework.
4.1.1 Matrix Factorization (MF)
Matrix Factorization is a foundational technique in recommendation systems and collaborative filtering. It decomposes a large, sparse interaction matrix (e.g., user-item ratings) into the product of two lower-rank latent factor matrices: $R\approx U^{\top}V,$ where $R$ is the observed rating matrix, $U$ and $V$ are matrices of user and item latent factors, respectively. The model is trained by minimizing reconstruction error while incorporating regularization to prevent overfitting:
$\min_{U,V}\sum_{(i,j)\in\mathcal{K}}(r_{ij}-U_{i}^{\top}V_{j})^{2}+\lambda(\|U\|^{2}+\|V\|^{2}),$
where $\mathcal{K}$ is the set of observed entries and $\lambda$ is a regularization parameter. The learned model $U^{\top}V$ can then be used to predict unknown ratings, effectively filling in the missing entries in the matrix (see Figure 1).
Figure 1: Illustration of the Matrix Factorization process.
In the SCF context, we can analogize users and items to entities such as stores and products, and ratings to observed sales. By identifying latent factors underlying these interactions, MF provides insight into complex, hidden relationships that simpler linear models may fail to capture.
4.1.2 Factorization Machines (FM)
While MF focuses on a user-item dyad, Factorization Machines extend this idea to a broader class of features and their interactions. FM efficiently captures pairwise feature interactions by introducing latent vectors for each feature:
$\hat{y}(\mathbf{x})=w_{0}+\sum_{i=1}^{n}w_{i}x_{i}+\sum_{i=1}^{n}\sum_{j=i+1}^{n}\langle\mathbf{v}_{i},\mathbf{v}_{j}\rangle x_{i}x_{j},$
[…]
Figure 2: From Matrix Factorization to Factorization Machines.
As shown in Figure 2, compared to MF, FM provides a more flexible structure that can incorporate a broader range of features, making it suitable for capturing complex relationships in diverse datasets. Additionally, compared to traditional polynomial kernel-based methods like Support Vector Machines (SVM), FM demonstrates superior generalization capabilities in sparse settings. The dense parametrization of SVMs relies on direct observations of feature interactions, which are often unavailable in sparse data, whereas FM can effectively estimate parameters even under such conditions (for more details, see Rendle (2010)).
4.1.3 Deep Factorization Machines (DeepFM)
Deep Factorization Machines combine the strengths of FM with deep neural networks to model both low-order and high-order feature interactions. The FM component efficiently learns second-order feature interactions, while the deep neural network component captures complex, higher-order interactions without the need for extensive manual feature engineering. The structure of DeepFM is shown in Figure 3.
Figure 3: Architecture of Deep Factorization Machines.
By integrating DeepFM, we can capture intricate relationships among various features of the covariates in SCF data (e.g., product attributes, store characteristics, advertisement factors), moving beyond linear approximations and paving the way for modeling complex sales distributions.
4.1.4 Quantile Regression Neural Network (QRNN)
The methods discussed earlier, including MF, FM, and DeepFM, are primarily designed for expectation-based prediction tasks. However, our goal is to predict the entire distribution of sales, rather than just the mean, which is crucial for robust SCF risk management. Given that DeepFM is essentially a neural network architecture, it becomes necessary to extend the linear quantile regression model of QRGMM into a neural network based quantile regression model. QRNN provides an ideal tool for achieving this extension.
QRNN enhances conventional quantile regression by leveraging neural networks to approximate quantile functions directly. Instead of constraining the quantile function to a linear form, QRNN employs a flexible neural network architecture—often composed of multiple dense layers with nonlinear activation functions—to model the potentially intricate relationship between covariates and various quantiles of the response variable. Unlike the loss functions used in MF, FM, or DeepFM—which typically minimize mean squared error—QRNN adopts the same pinball loss as the linear quantile regression model introduced earlier in this paper. In addition, QRNN provides an efficient way to fit multiple quantiles at different quantile levels $\tau_{j}$, the total loss is typically the sum over all quantiles and samples:
$\min_{\text{model parameters}}\sum_{j=1}^{m}\sum_{i=1}^{n}\rho_{\tau_{j}}(y_{i}-\hat{y}_{i,j}),$
where $\hat{y}_{i,j}$ denotes the predicted $\tau_{j}$-quantile for sample $i$, obtained as the output from the neural network model. The model can be efficiently solved by SGD methods.
By integrating this quantile-specific loss with a neural network structure, QRNN can learn multiple quantiles simultaneously, capturing distinct aspects of the underlying conditional distribution. Neural network components such as nonlinear activations, dropout layers, and proper regularization enable the QRNN to represent complex, non-linear patterns and maintain stability during training. In essence, QRNN extends quantile regression capabilities beyond linear assumptions, allowing for highly non-linear relationships and more realistic modeling of distributions.
4.2 Integration of DeepFM and QRGMM: QRGMM
Building on the techniques introduced above, we now present the integration of DeepFM and QRNN into the QRGMM framework, resulting in the QRGMM model. The idea is to leverage DeepFM’s ability to capture intricate feature interactions and QRNN’s flexibility in modeling quantiles to produce a generative model that can robustly simulate entire sales distributions under complex, realistic conditions. Its structure is shown in Figure 4.
Figure 4: Integration of DeepFM and QRNN (QRGMM).
By using DeepFM as the neural backbone of the quantile regression network, we ensure that the learned quantile functions are sensitive to intricate, second-order and higher-order feature interactions. Once the quantile functions are estimated, QRGMM’s inverse transformation approach is applied to generate samples. Specifically, for a given $\mathbf{x}$, we draw $u_{k}\sim\text{Uniform}(0,1)$, and interpolate between the estimated quantiles to generate synthetic observations $\hat{Y}_{k}(\mathbf{x})$ that approximate the conditional distribution of $Y(\mathbf{x})$. The overall procedure follows Algorithm 1, except that the quantile regression model is replaced with the QRGMM model based on DeepFM.
By integrating DeepFM and QRNN within the QRGMM framework, the resulting QRGMM model more accurately captures the complex, non-linear interactions among features that characterize SCF sales patterns, leading to more realistic sales distributions and better estimates of tail risks and other key distributional characteristics. This enhanced generative capability supports more informed lending decisions, improved estimation of risk measures, and stronger risk management strategies. By bridging the gap between linear QRGMM and real-world complexities, QRGMM significantly enhances the applicability of risk management techniques, enabling financial institutions to navigate the uncertainties of cross-border e-commerce markets with greater confidence and more effective risk mitigation.
5 Numerical Experiments
This section presents a comprehensive evaluation of our proposed QRGMM approach to assess its performance and practical value in SCF risk management. We begin with a synthetic scenario where the full data-generating process is known. In this controlled environment, we can thoroughly examine the model’s ability to accurately estimate conditional quantiles, generate realistic distributions, and precisely approximate various risk measures. Following the synthetic analysis, we apply the method to a real-world datasets, thereby demonstrating its relevance and adaptability to practical SCF contexts.
5.1 Experiments on Synthetic Data
To create a challenging yet interpretable test environment, we construct a synthetic dataset of 150,000 samples using an FM-Location-Scale Shift model:
$Y(\mathbf{x})=$ $w_{0}+\sum_{i=1}^{n}w_{i}x_{i}+\sum_{i=1}^{n}\sum_{j=i+1}^{n}\left\langle\mathbf{v}_{i},\mathbf{v}_{j}\right\rangle x_{i}x_{j}$
$+\left(r_{0}+\sum_{i=1}^{n}r_{i}x_{i}+\sum_{i=1}^{n}\sum_{j=i+1}^{n}\left\langle\mathbf{z}_{i},\mathbf{z}_{j}\right\rangle x_{i}x_{j}\right)u,$
where $\log(u)\sim N(0,1)$. Moreover, the covariate vector $\mathbf{x}=(x_{1},\ldots,x_{410})$ represents a mix of categorical and continuous features to mimic the complexity commonly encountered in SCF data. The first 100 features represent the one-hot encoding of product categories, the subsequent 300 features correspond to the one-hot encoding of seller IDs, and the final 10 features are continuous variables that capture additional product or market factors. This high-dimensional, heterogeneous covariate structure provides a realistic testbed for distribution-oriented modeling.
We randomly partition the data into a training set of 120,000 samples and a test set of 30,000 samples. The number of grid points for the model is set to $m=400$ in accordance with the recommendations provided in Hong et al. (2023) that optimal choice of $m=O(\sqrt{n})$ in practice. After the model is fitted, we assess its performance in terms of quantile estimation accuracy, generative quality, and risk measure estimation through subsequent evaluations, comparing it with the prior popular generative model, Conditional Wasserstein Generative Adversarial Networks with Gradient Penalty (CWGAN-GP, Athey et al. 2021, hereafter referred to as CWGAN).
5.1.1 Quantile Estimation and Distribution Generation
We first assess the accuracy of the model’s quantile estimation. For each test observation $\mathbf{x}_{i}$, we predict $\tau_{j}$-quantiles, denoted as $\hat{y}_{ij}(\tau)$, and compare the empirical proportion of test responses that fall below these predicted quantiles, $\hat{\tau}_{j}=\frac{1}{N_{\text{test}}}\sum_{i=1}^{N_{\text{test}}}\mathbb{I}(y_{i}\leq\hat{y}_{ij}(\tau))$, with the nominal level $\tau_{j}$, for $j=1,2,\cdots,m-1$. If the model perfectly estimates the $\tau_{j}$-quantile, $\hat{\tau}_{j}$ should be close to $\tau_{j}$, for $j=1,2,\cdots,m-1$. We conduct the experiment over 100 replications and plot the average $\hat{\tau}_{j}$ across these replications in Figure 5. The result demonstrates an excellent alignment between $\hat{\tau}_{j}$ and $\tau_{j}$ for all $j=1,2,\cdots,m-1$, confirming that the model provides accurate quantile estimates.
Figure 5: Alignment Between Nominal and Empirical Quantile Levels on the Synthetic Data.
Next, we evaluate the generative capabilities of the QRGMM model. To benchmark the generative performance, we conduct both unconditional and conditional tests, comparing the distributions of generated data by QRGMM and CWGAN model with the true test data.
Unconditional Test
Using the trained QRGMM, we generate synthetic observations $\hat{y}^{\prime}_{i}$ for each test covariate vector $\mathbf{x}_{i}^{\prime}$, thus constructing a synthetic test dataset. Then we can compare the overall distribution of the generated samples $\{\hat{y}^{\prime}_{i}:i=1,\ldots,N_{\text{test}}\}$ with the true test samples $\{y^{\prime}_{i}:i=1,\ldots,N_{\text{test}}\}$. We conducted 100 experimental replications, calculating the sample means and standard deviations for the generated datasets and true test dataset in each run. The averages of these values across the 100 replications are reported in Table 1. To quantify distributional similarity, we also computed the Wasserstein Distance (WD) and the Kolmogorov-Smirnov (KS) statistic between the generated datasets and the true test dataset for each replication. The histograms of these metrics across replications are illustrated in Figure 6.
Conditional Test
The above comparison is unconditional, assessing overall distributional similarity. Since we know the true data-generating process in the synthetic setting, we can also perform a conditional test to evaluate the generative performance conditioned on a specific covariate vector. For a given covariate vector $\mathbf{x}$, we generate conditional datasets of $\hat{Y}(\mathbf{x})$ using QRGMM and CWGAN model and compare them with the true conditional dataset of $Y(\mathbf{x})$ sampled from the true model. Like the unconditional test, we perform 100 replications, calculating the sample means and standard deviations for each conditional dataset. The averaged results are also included in Table 1. The WD and KS values for the conditional test across replications are shown in Figure 7.
Table 1: Mean and Standard Deviation (SD) of Generated Samples.
Model Unconditional Test Conditional Test
Mean SD Mean SD
Truth $4349.4819$ $1433.0318$ $3948.5480$ $148.1465$
QRGMM $4348.6113$ $1433.1995$ $3944.1098$ $144.8040$
CWGAN $4325.2962$ $1376.1644$ $3939.9981$ $187.7608$
Figure 6: WD (left) and KS (right) for Unconditional Test.
Figure 7: WD (left) and KS (right) for Conditional Test.
[…]
Figure 8: Histograms of generated vs. true data in one replication.
Additionally, we visualize the generative performance by presenting histograms of the generated and true data distributions from one replication in Figure 8 for both the unconditional test and conditional test.
Summary of Generative Performance
Our results demonstrate that QRGMM consistently achieves significantly lower WD and KS values compared to CWGAN in both unconditional and conditional tests (see Table 1 and Figures 67). This indicates that QRGMM provides a more accurate and robust reproduction of the true data-generating process, whereas CWGAN exhibits significant variability in performance across different replications. Furthermore, the histograms in Figure 8 illustrate that QRGMM-generated samples closely match the true data distribution in both unconditional and conditional settings, whereas CWGAN-generated samples exhibit noticeable discrepancies in conditional settings. These findings underscore the robustness and fidelity of QRGMM in modeling complex data distributions.
5.1.2 Risk Measures Approximation
In this synthetic experiment, for a specified $\mathbf{x}$, the data-generating process of $Y(\mathbf{x})$ is known, enabling us to analytically compute true risk measures as benchmarks for assessing the performance of QRGMM and CWGAN approximations. In our setup, the loss function $L(l)$ defined in equation (1) is used, and the maximum loan amount $\bar{l}$ is set to the 99th percentile of $rY(\mathbf{x})$, with net unit revenue $r=1$. Loan levels $l$ are evaluated at 100 equally spaced points across $[2500,\bar{l}]$.
We compute four true risk measures: $r_{1}(l)$ (probability of default), $r_{2}(l)$ (expected loss), and $\text{LGD}(l)$ (loss given default), as specified in Section 2, and $r_{3}(l)=\mathbb{E}[((l-rY(\mathbf{x}))^{+})^{2}]$, a specific form of the generalized risk measure highlighting severe losses. QRGMM and CWGAN approximations are calculated over 100 replications, each with 10,000 samples, and their means are compared to true values and estimations using the true test data samples averaged over 100 replications. These results are depicted in Figure 9, with subplots for $r_{1}(l)$, $r_{2}(l)$, $\text{LGD}(l)$, and $r_{3}(l)$.
Figure 9: Approximations of risk measures across loan levels $l$.
The results are highly encouraging: QRGMM’s approximations consistently track the true values across all measures, demonstrating robust accuracy. Conversely, CWGAN exhibits notable discrepancies, especially in $r_{2}(l)$ and $r_{3}(l)$, reflecting weaker performance. This comparison underscores QRGMM’s superior capability in capturing complex risk measures, QRGMM’s reliability significantly outshines CWGAN, demonstrating its applicability in SCF credit risk management. Notably, these figures vividly offer a clear, immediate view of how risk measures evolve with loan amount $l$, providing decision-makers with an intuitive tool for understanding the risk dynamics, thereby greatly facilitating loan sizing decisions.
5.2 Real-World Data Experiments
5.2.1 Big Mart Sales Dataset
To evaluate QRGMM in a practical setting, we apply it to the Big Mart Sales dataset (Shrivas 2013). This dataset contains 8,523 observations of 1,559 products across 10 stores, along with various product-level and store-level attributes. We randomly split the data into training and test subsets with a 4:1 ratio across 100 replications, and set $m=80$.
Quantile Estimation.
As in Section 5.1, we first examine the accuracy of the model’s quantile estimation. For each test observation, we estimate $\tau$-quantiles and compute the empirical proportion of actual sales below these predicted values, $\hat{\tau}$. Figure 10 shows that $\hat{\tau}$ closely aligns with $\tau$ across 100 replications, indicating that QRGMM continues to offer reliable quantile estimates on real-world data.
Figure 10: Alignment of Nominal and Empirical Quantile Levels on the Big Mart Data.
Unconditional Distribution Comparison.
Since the true data-generating mechanism is unknown in real-world datasets, we cannot perform a conditional test as in the synthetic experiment. Instead, we focus on an unconditional comparison of the generated and observed sales distributions. We trained the QRGMM and CWGAN models on the training set and subsequently generated a synthetic test dataset for each model across 100 replications.
Table 2 summarizes the means and standard deviations (averaged over the 100 replications) of both the generated and observed test sales. QRGMM shows a closer match to the true data than CWGAN, suggesting that QRGMM can capture the overall distribution of sales more accurately. To further quantify distributional similarity, we compute the WD and KS statistic for each replication, and plot their distributions in Figure 11. QRGMM again exhibits lower WD and KS values than CWGAN, reinforcing its superior generative ability in this real-world scenario.
Table 2: Mean and Standard Deviation of Generated Samples.
      Model       Mean       SD
      Truth       $2177.6046$       $1708.2316$
      QRGMM       $2168.4776$       $1680.5823$
      CWGAN       $2194.7852$       $1518.2655$
Figure 11: WD (left) and KS (right) for Big Mart Sales.
Figure 12 illustrates the observed and generated sales distributions in one representative replication. The QRGMM-generated distribution tracks the empirical distribution more closely than does CWGAN.
Figure 12: Histogram Comparison of Observed vs. Generated Big Mart Sales (Unconditional).
Discussion.
In summary, QRGMM continues to exhibit strong performance in quantile estimation and distribution generation on the Big Mart Sales dataset. Although a conditional test is infeasible due to unknown true mechanisms, the unconditional metrics and visual checks consistently show QRGMM outperforming CWGAN in capturing the real data distribution. This result underscores QRGMM’s practicality for complex SCF applications.
5.2.2 Amazon CBEC Dataset
To further assess the effectiveness of QRGMM in a real-world cross-border e-commerce (CBEC) context, we also conduct experiments on a dataset from a company conducting China-to-U.S. CBEC on Amazon. This dataset includes 3,495 sales records from the first quarter of 2023, covering 99 product categories across 131 stores. The products are managed by 23 distinct product selection teams and 23 sales teams, forming various team pairings. As noted earlier, these pairings exhibit combinatorial effects due to differences in team capability and cross-team collaboration. The dataset also features seven continuous variables: advertising expenses, clicks, impressions, sale price, product cost, average rating, and number of reviews. Again, we randomly split the data into training and test subsets with a 4:1 ratio across 100 replications. And we set $m=60$.
Quantile Estimation.
Similarly to Section 5.2.1, we begin by evaluating the model’s quantile estimation accuracy. For each test observation, we compute the empirical coverage $\hat{\tau}$ of the predicted $\tau$-quantiles across 100 replications. As shown in Figure 13, $\hat{\tau}$ aligns well with the nominal $\tau$, confirming that QRGMM maintains reliable quantile estimation on this CBEC dataset.
Figure 13: Alignment of Nominal and Empirical Quantile Levels on the Amazon CBEC Data.
Unconditional Distribution Comparison.
Since the true conditional distribution remains unknown, we again adopt an unconditional evaluation strategy. Both QRGMM and CWGAN are trained on the training set and used to generate synthetic test sales data over 100 replications.
Table 3 reports the means and standard deviations (averaged across 100 replications) of the generated and observed test sales. Similar to the results on Big Mart dataset, QRGMM aligns more closely with the true data than CWGAN. In addition, Figure 14 compares the WD and KS statistics across replications. QRGMM consistently yields lower values for both metrics, demonstrating its superior generative fidelity in this CBEC scenario.
Table 3: Mean and Standard Deviation of Generated Samples.
      Model       Mean       SD
      Truth       $44.2296$       $102.2226$
      QRGMM       $39.2979$       $76.7818$
      CWGAN       $36.0173$       $60.7141$
Figure 14: WD (left) and KS (right) for Amazon CBEC Sales.
Figure 15 presents a histogram comparison of observed and generated sales distributions from one representative replication. As with the Big Mart dataset, the QRGMM-generated samples align more closely with the observed sales, whereas CWGAN shows noticeable deviation.
Figure 15: Histogram Comparison of Observed vs. Generated Amazon CBEC Sales (Unconditional).
Discussion.
Overall, QRGMM continues to demonstrate robust performance in both quantile estimation and unconditional distribution generation on the Amazon CBEC dataset. Despite the complex interaction effects among product and sales teams, as well as the influence of various marketing and product attributes, the model effectively captures the data structure. These results affirm QRGMM’s practical utility in complex CBEC environments, enhancing its relevance for SCF risk management.
6 Conclusion
This paper introduced a novel SCF risk management framework that leverages generative model to provide a full distributional understanding of sales, developing a unified analytical framework that simultaneously addresses credit risk assessment and loan size determination—objectives traditionally treated separately. By integrating QRGMM—a quantile-regression-based generative modeling method—with DeepFM to capture complex feature interactions, we have developed a flexible approach capable of estimating entire conditional distributions in complex SCF settings. This allows practitioners to move beyond conventional point estimations and simple risk classifications, offering a data-driven, distribution-oriented perspective on SCF risk.
Building on the linear quantile regression framework of Hong et al. (2023), we establish rigorous asymptotic guarantees for three key risk measures, which serve as universal building blocks for deriving a wide range of risk metrics. Numerical experiments on synthetic data confirm that QRGMM accurately and stably approximates the underlying distribution, surpasses CWGAN in unconditional and conditional distribution matching, and recovers risk measures that align closely with known truths across the range of loan level space, greatly facilitating loan-sizing decision making in an intuitive way. On real-world datasets, QRGMM demonstrates practical applicability, accurately estimating quantiles and generating distributions that mirror real sales patterns.
In summary, this work represents an initial exploration of the use of generative AI for operations management, advancing existing SCF risk management methodologies by employing generative models that provide richer distributional insights. However, much work remains to be done in this area. Future research should extend the framework to handle time-series data with seasonal effects and tackle challenges related to high-dimensional datasets. Additionally, while this study focuses on a single product for an SME, future efforts could explore loan portfolio problems by accounting for intricate interdependencies among products, including complementarity and substitution effects. And applications in other domains such as financial risk management and supply chain optimization could also be investigated. By continuing to refine these generative distributional approaches, the ultimate goal is to support more informed, resilient, and data-driven decision-making in various fields of operations management.